Determine the expected count for each outcome.

N=633

i

1

2

3

4

pi

0.16

0.35

0.16

0.33

The expected count for the outcome 1 ___________.(Round to two decimal places as needed).

The expected count for outcome 2 is____________.(Round to two decimal places as needed).

The expected count for outcome 3 is___________.(Round to two decimal places as needed)

The expected count for outcome 4 is___________(Round to two decimal places as needed)
Answer #1

The expected count for the outcome 1 is = 633 ( 0.16 )

= 101.28

The expected count for the outcome 2 is = 633 ( 0.35 )

= 221.55

The expected count for the outcome 3 is = 633 ( 0.16 )

= 101.28

The expected count for the outcome 4 is = 633 ( 0.33 )

= 208.89